The line through the points (m, -9) and (7, m) has slope m. What is the value of m?
A. 3 B. -7/9 C. 16 D. 5 E. None of these

Answer by richwmiller(17219)

(-9-m)/(m-7)=m
-9-m=m^2-7m
0=m^2-6m+9
0=(m-3)^2
m=3

substituting 3 for m
-9-3/3-7=3
-12/-4=3
3=3
ok